OpenStudy (anonymous):

what is the inverse function of f(x)=(e^x)/1+8e^x

OpenStudy (dumbcow):

flip x and y \[x = \frac{e^y}{1+8e^y}\] \[e^y = x + 8x e^y\] \[e^y (1-8x) = x\] \[e^y = \frac{x}{1-8x}\] \[y = \ln (\frac{x}{1-8x})\] For x>0 , x not equal 1/8